Consider a bond with a

77​%

annual coupon and a face value of

​$1100.

Complete the following table. ​(Enter your responses rounded to two decimal​ places.)

years to maturity yield to maturity current price

2 5%

2 7%

3 7%

5 5%

5 9%
Answer #1

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = __% (YTM of bond)

nper = __ (__ years remaining until maturity with 1 coupon payment each year)

pmt = 1100 * 7.7% (annual coupon payment = face value * coupon rate)

fv = 1100 (face value receivable on maturity)

A B Years to 1 maturity YTM 5% 7% 7% Current price ($1,155.22) ($1,113.92) ($1,120.21) ($1,228.59) ($1,044.38) u uw 5% 9%

- AB Years to 1 maturity YTM 22 0.05 3 2 0.07 43 0.07 5 5 0.05 0.09 Current price =PV(B2, A2, 1100*7.7%, 1100) =PV(B3,A3,1100